OpenStudy (anonymous):

The graph of a function is horizontally compressed by a factor of 5 and vertically compressed by a factor of 2 Find an equation for this compressed function in terms of the function f(x)

OpenStudy (whpalmer4):

Horizontal compression means that the value of x given to the function grows faster than x. So, if your function = 0 at x = 0, and 4 at x = 2, compressing it by a factor of 2 horizontally would make the function = 0 at x = 0, and 4 at x = 1. A little thought should show that this is the same as writing f(2*x) or f(2x). Now how about the vertical compression? Do we want to multiply, or divide, by our scale factor?

OpenStudy (anonymous):

i thought the answer was y=2f(5x) but it comes back as wrong im not sure whats is wrong. i know the a=vertical b=horizontal

OpenStudy (whpalmer4):

Ah, but you're *compressing* the vertical scale, right? Is that multiplying by 2, or dividing by 2?

OpenStudy (anonymous):

divide?

OpenStudy (whpalmer4):

You tell me. Say your function is f(x) = x^2. x y = x^2 y = 2x^2 y = 1/2 x^2 0 0 0 0 1 1 2 1/2 2 4 8 2 3 9 18 4 1/2 which one of those looks like it is y = x^2 vertically compressed by a factor of 2?

OpenStudy (anonymous):

the last one

OpenStudy (whpalmer4):

Right, for vertical compression, we divide by the compression factor. So your answer to the problem is?

OpenStudy (anonymous):

1/2f(5x)